OpenStudy (anonymous):

Write as a single log... 3lna-1/4(3lnb+lnc) *ln is natural log. a, b, and c are variables

OpenStudy (anonymous):

\[3lna =lna^3 .......(i) \& \frac{1}{4}(3lnb+lnc) = \frac{1}{4}lnb^3c = \ln(b^3c)^{\frac{1}{4}} (ii) using (i) and (ii) u get lna^3-\ln(b^3c)\]